OpenStudy (scorcher219396):

Find the derivative of 1/(sqrt(x+2)) using the definition of a derivative

OpenStudy (scorcher219396):

Work so far:

OpenStudy (scorcher219396):

I'm just not sure what to do next?

OpenStudy (scorcher219396):

@Kainui

OpenStudy (kainui):

Ahh the trick is to multiply by the conjugate. All that means is you make a fraction that looks like the two square roots on top but with a different sign inbetween. \[\Large \frac{\sqrt{a}+\sqrt{b}}{\sqrt{a}+\sqrt{b}}\] That will get rid of the square roots on the top allowing you to get rid of your h on the bottom.

OpenStudy (scorcher219396):

So i'm aiming to get rid of the numerator of the numerator? (the top line?) I was assuming I would need to rationalize the 2nd line

OpenStudy (kainui):

The main thing you're trying to get rid of is the h in the denominator because that is essentially dividing by zero and that's what's preventing you from plugging in h=0. So if you can get an h to show up in the numerator, then it will divide away and you can plug it in safely.

OpenStudy (scorcher219396):

Ok thanks, that makes a lot of sense